# A Free Probabilistic Framework for Denoising Diffusion Models: Entropy, Transport, and Reverse Processes

### v1

@@ Abstract
This work develops a rigorous framework for diffusion-based generative modeling in the setting of free probability. We extend classical denoising diffusion probabilistic models to free diffusion processes—stochastic dynamics acting on noncommutative random variables whose spectral measures evolve by free additive convolution. The forward dynamics satisfy a free Fokker–Planck equation that increases Voiculescu’s free entropy and dissipates free Fisher information, providing a noncommutative analogue of the classical de Bruijn identity. Using tools from free stochastic analysis, including a free Malliavin calculus and a Clark–Ocone representation, we derive the reverse-time stochastic differential equation driven by the conjugate variable, the free analogue of the score function. We further develop a variational formulation of these flows in the free Wasserstein space, showing that the resulting gradient-flow structure converges to the semicircular equilibrium law. Together, these results connect modern diffusion models with the information geometry of free entropy and establish a mathematical foundation for generative modeling with operator-valued or high-dimensional structured data.
@@ I Introduction
However, the existing theory remains fundamentally commutative. The noise is Gaussian, the random variables inhabit Euclidean spaces, and correlations are captured by classical covariance. In contrast, contemporary machine learning increasingly operates in operator-valued and inherently non-commutative regimes. Large neural networks, for example, are more naturally characterized by their random matrix spectra rather than by vector-valued features, as demonstrated in empirical spectral analyses of deep architectures [8, 9, 10]. In such settings, classical probability fails to capture the operator-level correlations and non-commuting dependencies that emerge from matrix multiplications and compositional network structure. This gap motivates a search for a non-commutative foundation for generative modeling.
Free probability theory provides precisely such a foundation. Introduced by Voiculescu [11, 12] and developed further by Biane [13] and Biane–Speicher [14], free probability extends stochastic calculus, entropy, and information theory to non-commutative random variables. Its connections to random matrix theory are well established: empirical spectral distributions converge to free convolutions, and asymptotic freeness governs large-dimensional limits [15]. Moreover, the theory introduces free analogues of fundamental information measures—free entropy and free Fisher information—which mirror their Shannon counterparts [12]. The development of free optimal transport by Biane and Voiculescu [16] and its extensions in [17] established a non-commutative analogue of the $L_{2}$–Wasserstein space. More recently, free probabilistic ideas have begun to influence deep learning theory itself: for example, Das [18] used operator-valued free convolutions to analyze Transformer attention as a form of non-commutative stochastic interaction.
These developments naturally raise the following question: Can diffusion-based generative modeling be extended into the free, non-commutative domain, while retaining the geometric and information-theoretic structure of classical diffusions?
This paper provides an affirmative answer. We develop a rigorous framework for free diffusion models, extending denoising diffusion processes into the realm of free probability. By formulating both the forward and reverse dynamics in terms of free stochastic calculus, we establish explicit connections between free entropy, free Fisher information, and non-commutative Wasserstein geometry. Our analysis further leads to a family of functional inequalities—including free analogues of the logarithmic Sobolev, Talagrand, and HWI inequalities (linking H for relative entropy, W for Wasserstein distance, and I for Fisher information)—that govern convergence, stability, and transport in the free setting.
@@
  • Information-theoretic and variational structure: We prove a free de Bruijn identity linking entropy production to free Fisher information, show that the forward process forms a gradient ascent of free entropy $\chi$, and construct a JKO-type variational scheme in the free Wasserstein metric.
@@
  • Free Schrödinger bridge and energy–entropy balance: Through a free Girsanov theorem, we characterize the reverse diffusion as a free Schrödinger bridge, revealing a precise energy–entropy trade-off in the non-commutative setting.
@@
Taken together, these results position free diffusion models as Wasserstein gradient flows of Voiculescu’s free entropy, unifying stochastic diffusion, functional inequalities, and non-commutative information geometry under a single rigorous framework. Beyond serving as a conceptual extension of classical diffusion models, this theory introduces new analytical tools for operator-valued and high-dimensional structured data, opening a research program at the intersection of free probability, optimal transport, and information-theoretic generative modeling.
@@ II-C Information-Theoretic Viewpoint
This duality between entropy production (forward) and entropy dissipation (reverse) provides a geometric and information-theoretic foundation for diffusion-based generative modeling. It also sets the stage for the non-commutative generalization developed later in this paper, where Shannon entropy and Fisher information are replaced by their free-probabilistic counterparts.
@@ II-D Applications and Impact
Diffusion models have transformed several areas of artificial intelligence:
  • Image, audio, and video generation: Samples from the reverse SDE (4) achieve state-of-the-art realism in large-scale models such as Imagen, DALL-E 2, and Stable Diffusion [24, 25, 26].
  • Conditional generation: Conditioning mechanisms, such as classifier guidance, enable controllable text-to-image or text-to-audio synthesis.
  • Uncertainty quantification: Score-based methods provide flexible tools for approximate Bayesian inference and posterior sampling [4].
  • Scientific and physical modeling: Diffusion-based SDEs have been used to generate molecular conformations, simulate turbulent flows, and model protein structures.
From a theoretical standpoint, the universality of diffusion models lies in their two complementary limits: the forward diffusion maps any data distribution to a Gaussian prior, while the reverse process—given an accurate score estimate—can recover arbitrarily complex distributions. This classical framework will serve as the benchmark for the free diffusion models developed in subsequent sections, where we extend these principles to non-commutative, operator-valued random variables.
@@ III Free Probability Formulation
Classical diffusion-based generative models operate in Euclidean probability spaces, where random variables commute and independence is described by factorization of expectations. However, in many high-dimensional or structured learning systems—such as deep neural networks, random feature maps, or transformer attention operators—the effective random objects are not scalars but matrices or operators. Their entries are highly correlated, yet their large-dimensional behavior often exhibits a form of asymptotic freeness rather than classical independence. This observation motivates the use of free probability theory as a natural non-commutative extension of diffusion modeling.
@@ III-A Free Probability Background
This spectral law generalizes the notion of a probability density to the non-commutative world.
@@ III-B Freeness
whenever $a_{j}\in\mathcal{A}_{i_{j}}$, $\varphi(a_{j})=0$, and consecutive indices differ ($i_{1}\neq i_{2}\neq\cdots\neq i_{k}$). This property means that alternating, centered elements from distinct subalgebras have zero expectation—the non-commutative analogue of the classical product rule $\mathbb{E}[XY]=\mathbb{E}[X]\mathbb{E}[Y]$ for independent variables.
@@ III-D Free Diffusion as a Model of Operator-Valued Noise
Consider now a non-commutative diffusion process
@@
where $X_{t},b_{t},\sigma_{t}\in\mathcal{A}_{\mathrm{sa}}$ and $S_{t}$ is free Brownian motion. The spectral measure $\mu_{X_{t}}$ evolves by free additive convolution—a non-commutative analogue of Gaussian convolution. When $b_{t}=-\tfrac{1}{2}\beta(t)X_{t}$ and $\sigma_{t}=\sqrt{\beta(t)}$, this becomes a free Ornstein–Uhlenbeck process [19], whose stationary law is the semicircular distribution.
@@ III-F Summary
Free probability thus provides a rigorous language for describing diffusion in non-commutative or matrix-valued settings. The forward process corresponds to the flow of spectral measures under free additive convolution, while the reverse process represents an entropy-gradient descent in the geometry of free Wasserstein space. This framework extends the statistical mechanics interpretation of DDPMs to the realm of operator algebras, offering a foundation for analyzing and potentially designing generative models for structured, non-commuting data.
@@ V Free Fokker–Planck Dynamics and Free Clark–Ocone Formula
Once the forward diffusion is lifted from classical probability to a non-commutative setting, the law of a free diffusion process $X_{s}$ evolves according to a free Fokker–Planck equation. Let $\mu_{X_{s}}$ denote the spectral distribution of $X_{s}$ in a tracial $W^{*}$-probability space $(\mathcal{A},\varphi)$, with dynamics
where $(S_{s})_{s\geq 0}$ is a free Brownian motion [13]. Then $\mu_{X_{s}}$ satisfies, in the distributional sense,
$\frac{\partial}{\partial s}\mu_{X_{s}}(x)=\frac{1}{2}\beta(s)\frac{\partial}{\partial x}\big(x\mu_{X_{s}}(x)\big)+\frac{1}{2}\beta(s)\frac{\partial^{2}}{\partial x^{2}}\mu_{X_{s}}(x),$ (8)
which is a free counterpart of the classical Fokker–Planck equation. Equation (8) governs the evolution of the spectral law, and can be shown to correspond to a gradient flow of Voiculescu’s free entropy $\chi(\mu_{X_{s}})$ in the metric space of spectral distributions endowed with the free 2-Wasserstein distance [14, 27].
@@
Let $(S_{t})_{t\geq 0}$ be a free Brownian motion on $(\mathcal{A},\varphi)$, and let $F\in L^{2}(\mathcal{A},\varphi)$ be square-integrable with respect to the non-commutative filtration $(\mathcal{F}_{t})$ generated by $S_{t}$. Then there exists a unique adapted process $\mathbb{E}[\mathcal{D}_{u}F\mid\mathcal{F}_{u}]$ such that
$F=\mathbb{E}[F]+\int_{0}^{T}\mathbb{E}[\mathcal{D}_{u}F\mid\mathcal{F}_{u}]\,\mathrm{d}S_{u},$ (9)
where $\mathcal{D}_{u}$ denotes the free Malliavin derivative. This decomposition is the non-commutative analogue of the classical Clark–Ocone representation [14, 28].
@@ VI Reverse-Time Free SDE: Malliavin and PDE Derivations
extends Voiculescu’s classical free integration-by-parts identity from non-commutative polynomials to the full Sobolev space $W^{1,2}_{\mathrm{free}}(\mu_{t})$.
@@
Let $F\in\mathrm{Dom}(D)$ and approximate it by a sequence of self-adjoint non-commutative polynomials $\{P_{n}(X_{t})\}_{n\geq 1}$ such that $P_{n}(X_{t})\to F$ in $L^{2}(\mathcal{A},\tau)$ and $DP_{n}(X_{t})\to DF$ in $L^{2}(\mathcal{A}\otimes\mathcal{A})$. For each polynomial $P_{n}$, Voiculescu’s original integration-by-parts formula [12, Prop. 7.1] gives
@@
where $\mu_{X_{0}}$ is compactly supported with $\chi(\mu_{X_{0}})>$-$\infty$ and $\Phi^{\ast}(\mu_{X_{0}})<\infty$. Assume that $(X_{t})_{0\leq t\leq T}$ exists uniquely in the free Itô sense, and that each $\mu_{X_{t}}$ admits a conjugate variable $\Xi_{t}\in L^{2}(W^{\ast}(X_{t}),\varphi)$. Then the time-reversed process $\widetilde{X}_{s}\mathrel{\mathop{\ordinarycolon}}=X_{T-s}$ satisfies the stochastic differential equation
@@
where $(\widetilde{S}_{s})_{s\in[0,T]}$ is a backward free Brownian motion adapted to the reversed filtration $\widetilde{\mathcal{F}}_{s}=\sigma\{\widetilde{X}_{u}\mathrel{\mathop{\ordinarycolon}}0\leq u\leq s\}$.
@@
Step 1: Existence and forward representation. Let $\Lambda(t)=\int_{0}^{t}\beta(r)\,\mathrm{d}r$. By free Itô integration theory [14], the SDE (F) admits the explicit solution
$X_{t}=e^{-\frac{1}{2}\Lambda(t)}X_{0}+\!\int_{0}^{t}e^{-\frac{1}{2}(\Lambda(t)-\Lambda(u))}\sqrt{\beta(u)}\,\mathrm{d}S_{u}.$
The free Itô isometry ensures that $\|X_{t}\|_{2}^{2}<\infty$ and thus $X_{t}\in L^{2}(\mathcal{A},\varphi)$ for all $t\in[0,T]$.
Step 2: Clark–Ocone representation. For each fixed $T$, the free Clark–Ocone formula (see Appendix C for a rigorous derivation) gives an adapted representation of $X_{T}$:
@@
Substituting the explicit solution of $X_{T}$ into the stochastic integral and applying the free Itô isometry yields
@@
This follows because $(\mathrm{d}S_{t})_{t\geq 0}$ has freely independent increments and $u_{t}$ is adapted; hence the conditional expectation $\mathbb{E}_{\varphi}[\cdot\mid\mathcal{F}_{t}]$ eliminates all stochastic terms beyond time $t$, leaving only the deterministic damping factor $e^{-\frac{1}{2}(\Lambda(T)-\Lambda(t))}$.
Step 3: Identification of the drift correction. By Lemma 3, for the log-density functional $G=\log\pi_{t}(X_{t})$ one has $\mathbb{E}_{\varphi}[D_{t}G\mid\mathcal{F}_{t}]=-\,\Xi_{t}$. Combining this with the derivative term $\tfrac{1}{2}\beta(t)X_{t}$ in the forward SDE (F) gives the backward drift $\tfrac{1}{2}\beta(t)X_{t}-\beta(t)\Xi_{t}$.
Step 4: Time reversal. Define $\widetilde{X}_{s}=X_{T-s}$. By free stochastic calculus with respect to backward filtrations (see Biane [29], Sec. 3), the differential of $\widetilde{X}_{s}$ satisfies
where $\widetilde{S}_{s}\mathrel{\mathop{\ordinarycolon}}=S_{T}-S_{T-s}$ is a free Brownian motion under the reversed filtration. The integrability of the drift and diffusion coefficients follows from the square-integrability of $X_{t}$ and $\Xi_{t}$, ensuring well-posedness of (R) in $L^{2}(\mathcal{A},\varphi)$.
Hence, the time-reversed process $\widetilde{X}_{s}$ is a weak solution of (R), completing the proof. ∎
@@
Both derivations show that the term $\Xi_{s}$ plays the role of the non-commutative score function in the reverse-time dynamics, and that the reversed diffusion acts as a free deconvolution flow.

### v2

@@ Abstract
This paper develops a rigorous probabilistic framework that extends denoising diffusion models to the setting of noncommutative random variables. Building on Voiculescu’s theory of free entropy and free Fisher information, we formulate diffusion and reverse processes governed by operator-valued stochastic dynamics whose spectral measures evolve by additive convolution. Using tools from free stochastic analysis—including a Malliavin calculus and a Clark–Ocone representation—we derive the reverse-time stochastic differential equation driven by the conjugate variable, the analogue of the classical score function. The resulting dynamics admit a gradient-flow structure in the noncommutative Wasserstein space, establishing an information-geometric link between entropy production, transport, and deconvolution. We further construct a variational scheme analogous to the Jordan–Kinderlehrer–Otto (JKO) formulation and prove convergence toward the semicircular equilibrium. The framework provides functional inequalities (free logarithmic Sobolev, Talagrand, and HWI) that quantify entropy dissipation and Wasserstein contraction. These results unify diffusion-based generative modeling with the geometry of operator-valued information, offering a mathematical foundation for generative learning on structured and high-dimensional data.
@@ I Introduction
However, the existing theory remains fundamentally commutative. The noise is Gaussian, the random variables inhabit Euclidean spaces, and correlations are captured by classical covariance. In contrast, contemporary machine learning increasingly operates in operator-valued and inherently noncommutative regimes. Large neural networks, for example, are more naturally characterized by their random matrix spectra rather than by vector-valued features, as demonstrated in empirical spectral analyses of deep architectures [8, 9, 10]. In such settings, classical probability fails to capture the operator-level correlations and non-commuting dependencies that emerge from matrix multiplications and compositional network structure. This gap motivates a search for a noncommutative foundation for generative modeling.
Free probability theory provides precisely such a foundation. Introduced by Voiculescu [11, 12] and developed further by Biane [13] and Biane–Speicher [14], free probability extends stochastic calculus, entropy, and information theory to noncommutative random variables. Its connections to random matrix theory are well established: empirical spectral distributions converge to free convolutions, and asymptotic freeness governs large-dimensional limits [15]. Moreover, the theory introduces free analogues of fundamental information measures—free entropy and free Fisher information—which mirror their Shannon counterparts [12]. The development of free optimal transport by Biane and Voiculescu [16] and its extensions in [17] established a noncommutative analogue of the $L_{2}$–Wasserstein space. More recently, free probabilistic ideas have begun to influence deep learning theory itself: for example, Das [18] used operator-valued free convolutions to analyze Transformer attention as a form of noncommutative stochastic interaction.
These developments naturally raise the following question: Can diffusion-based generative modeling be extended into the free, noncommutative domain, while retaining the geometric and information-theoretic structure of classical diffusions?
This paper provides an affirmative answer. We develop a rigorous framework for free diffusion models, extending denoising diffusion processes into the realm of free probability. By formulating both the forward and reverse dynamics in terms of free stochastic calculus, we establish explicit connections between free entropy, free Fisher information, and noncommutative Wasserstein geometry. Our analysis further leads to a family of functional inequalities—including free analogues of the logarithmic Sobolev, Talagrand, and HWI inequalities (linking H for relative entropy, W for Wasserstein distance, and I for Fisher information)—that govern convergence, stability, and transport in the free setting.
I-A Outline and Contributions
@@
  • Information-theoretic and variational structure: We prove a free de Bruijn identity linking entropy production to free Fisher information, show that the forward process forms a gradient ascent of free entropy $\chi$, and construct a JKO (Jordan–Kinderlehrer–Otto)-type variational scheme in the free Wasserstein metric.
@@
  • Free Schrödinger bridge and energy–entropy balance: Through a free Girsanov theorem, we characterize the reverse diffusion as a free Schrödinger bridge, revealing a precise energy–entropy trade-off in the noncommutative setting.
@@
I-B Information-Theoretic Implications
The correspondence between free and classical entropy flows reveals a deep information-theoretic parallel. In classical diffusion-based generative models, the Fokker–Planck evolution drives Shannon entropy increase, and the reverse-time process recovers structure by following the gradient of the log-likelihood—a manifestation of entropy dissipation balanced by Fisher information flow. In the free setting, Voiculescu’s free entropy $\chi(\mu_{t})$ obeys an analogous dissipation law $\frac{\mathrm{d}}{\mathrm{d}t}\chi(\mu_{t})=\tfrac{1}{2}\Phi^{\ast}(\mu_{t})$, where $\Phi^{\ast}$ is the free Fisher information. This duality mirrors the de Bruijn identity and encodes how spectral measures lose and regain information under noncommutative diffusion. The reverse free SDE thus performs an information-preserving transport analogous to generative decoding, while the forward flow acts as a compression map toward a semicircular prior. Consequently, free diffusion models naturally admit a rate–distortion interpretation: entropy decay quantifies compression of operator-valued data, and the reverse flow reconstructs high-entropy signals along optimal information-geometric trajectories.
Taken together, these results position free diffusion models as Wasserstein gradient flows of Voiculescu’s free entropy, unifying stochastic diffusion, functional inequalities, and noncommutative information geometry under a single rigorous framework. Beyond serving as a conceptual extension of classical diffusion models, this theory introduces new analytical tools for operator-valued and high-dimensional structured data, opening a research program at the intersection of free probability, optimal transport, and information-theoretic generative modeling.
@@ II-C Information-Theoretic Viewpoint
This duality between entropy production (forward) and entropy dissipation (reverse) provides a geometric and information-theoretic foundation for diffusion-based generative modeling. It also sets the stage for the noncommutative generalization developed later in this paper, where Shannon entropy and Fisher information are replaced by their free-probabilistic counterparts.
@@ II-D Applications and Impact
Diffusion models have rapidly emerged as one of the most influential frameworks in modern artificial intelligence, reshaping both the theoretical and practical landscape of generative modeling. Their success stems from a simple but powerful principle: by learning to reverse a gradual corruption process that transforms structured data into noise, these models can synthesize entirely new samples that are statistically indistinguishable from real data. Recent surveys such as [7] have highlighted their broad theoretical foundations and the unifying stochastic framework they provide across disciplines.
  • High-fidelity image, audio, and video generation: Diffusion-based architectures now dominate large-scale generative modeling. Systems such as Imagen, Stable Diffusion, and DALL–E 2 produce photorealistic images, audio signals, and video clips through the reverse stochastic differential equation (SDE) (4). The ability to traverse from Gaussian noise to rich data manifolds via controlled stochastic dynamics has become a cornerstone of creative AI, surpassing the fidelity of GAN and flow-based models [24, 25, 26].
  • Conditional and controllable synthesis: By integrating classifier guidance, cross-modal conditioning, or attention-based mechanisms, diffusion models enable text-to-image, image-to-speech, and even multimodal generation. The probabilistic structure of these models naturally accommodates conditioning and interpolation, providing a flexible platform for controllable synthesis and style transfer.
  • Bayesian inference and uncertainty quantification: The connection between diffusion models and stochastic control [27, 4] allows them to perform approximate Bayesian inference. Score-based sampling schemes provide uncertainty estimates and enable posterior sampling in inverse problems, leading to applications in reinforcement learning, climate modeling, and medical imaging.
  • Scientific and physical simulation: Diffusion-based stochastic dynamics have been successfully applied to model complex physical phenomena, ranging from molecular conformations and quantum energy landscapes to turbulent and fluid flows. Recent works have demonstrated diffusion generative models for molecular structure prediction [28, 29] and high-energy physics through calorimeter simulation frameworks [30]. In these cases, diffusion models act as statistical surrogates for high-dimensional dynamical systems, blending physics-informed priors with generative flexibility.
Theoretical significance.
From a mathematical standpoint, diffusion models exhibit a remarkable universality: the forward diffusion maps any data distribution to a simple Gaussian reference, while the reverse-time SDE—parameterized by the score function—can, in principle, recover any target distribution. This bidirectional structure links the theory of diffusion models to the geometry of the Wasserstein space and to entropy-regularized optimal transport [31, 32]. It also provides a continuous-time view of deep generative learning that bridges stochastic control, variational inference, and information geometry.
Motivation for the free probabilistic framework.
However, many emerging machine learning settings involve data that are inherently noncommutative or operator-valued. Examples include covariance matrices in signal processing, weight matrices in deep networks, density operators in quantum information, and graph Laplacians in network analysis. In such regimes, classical diffusion models—built on Euclidean and commutative probability—cannot capture the higher-order spectral interactions or non-commuting dependencies that drive these systems.
The free probabilistic framework developed in this paper extends diffusion modeling into this operator-valued domain. By replacing Gaussian noise with semicircular noise and classical entropy with free entropy, we obtain a noncommutative analogue of the diffusion process that evolves along the geometry of the free Wasserstein space [16, 17]. Practitioners can view this as a theoretical foundation for generative modeling of structured, matrix- or tensor-valued data—offering new analytical tools for understanding generalization, stability, and the spectral behavior of large models.
Ultimately, free diffusion models bridge stochastic generative modeling with the deep structure of random matrix theory and noncommutative information geometry, suggesting a unified perspective on data generation across classical and operator-valued domains.
@@ III Free Probability Formulation
Classical diffusion-based generative models operate in Euclidean probability spaces, where random variables commute and independence is described by factorization of expectations. However, in many high-dimensional or structured learning systems—such as deep neural networks, random feature maps, or transformer attention operators—the effective random objects are not scalars but matrices or operators. Their entries are highly correlated, yet their large-dimensional behavior often exhibits a form of asymptotic freeness rather than classical independence. This observation motivates the use of free probability theory as a natural noncommutative extension of diffusion modeling.
@@ III-A Free Probability Background
This spectral law generalizes the notion of a probability density to the noncommutative world.
@@ III-B Freeness
whenever $a_{j}\in\mathcal{A}_{i_{j}}$, $\varphi(a_{j})=0$, and consecutive indices differ ($i_{1}\neq i_{2}\neq\cdots\neq i_{k}$). This property means that alternating, centered elements from distinct subalgebras have zero expectation—the noncommutative analogue of the classical product rule $\mathbb{E}[XY]=\mathbb{E}[X]\mathbb{E}[Y]$ for independent variables.
@@ III-D Free Diffusion as a Model of Operator-Valued Noise
Consider now a noncommutative diffusion process
@@
where $X_{t},b_{t},\sigma_{t}\in\mathcal{A}_{\mathrm{sa}}$ and $S_{t}$ is free Brownian motion. The spectral measure $\mu_{X_{t}}$ evolves by free additive convolution—a noncommutative analogue of Gaussian convolution. When $b_{t}=-\tfrac{1}{2}\beta(t)X_{t}$ and $\sigma_{t}=\sqrt{\beta(t)}$, this becomes a free Ornstein–Uhlenbeck process [19], whose stationary law is the semicircular distribution.
@@ III-F Summary
Free probability thus provides a rigorous language for describing diffusion in noncommutative or matrix-valued settings. The forward process corresponds to the flow of spectral measures under free additive convolution, while the reverse process represents an entropy-gradient descent in the geometry of free Wasserstein space. This framework extends the statistical mechanics interpretation of DDPMs to the realm of operator algebras, offering a foundation for analyzing and potentially designing generative models for structured, non-commuting data.
@@ V Free Fokker–Planck Dynamics and Free Clark–Ocone Formula
Once the forward diffusion is lifted from classical to noncommutative probability, the law of a free diffusion process $X_{s}$ evolves according to a free Fokker–Planck equation. Let $\mu_{X_{s}}$ denote the spectral measure of $X_{s}$ in a tracial $W^{*}$-probability space $(\mathcal{A},\phi)$ satisfying
$dX_{s}=-\tfrac{1}{2}\beta(s)X_{s}\,ds+\sqrt{\beta(s)}\,dS_{s},$
where $(S_{s})_{s\geq 0}$ is a free Brownian motion. Then the family of spectral measures $\{\mu_{X_{s}}\}_{s\geq 0}$ obeys, in the distributional sense,
$\frac{\partial\mu_{X_{s}}}{\partial s}(x)=\frac{1}{2}\beta(s)\frac{\partial}{\partial x}\!\left[x\,\mu_{X_{s}}(x)\right]+\frac{1}{2}\beta(s)\frac{\partial^{2}\mu_{X_{s}}}{\partial x^{2}}(x),$ (8)
which represents the commutative projection of the noncommutative evolution. Equation (8) is the gradient flow of Voiculescu’s free entropy $\chi(\mu_{X_{s}})$ in the metric space $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$ introduced by Biane and Voiculescu [33]. Equation (8) governs the evolution of the spectral law, and can be shown to correspond to a gradient flow of Voiculescu’s free entropy $\chi(\mu_{X_{s}})$ in the metric space of spectral distributions endowed with the free $L_{2}$-Wasserstein distance [14, 34].
@@
Let $(S_{t})_{t\geq 0}$ be a free Brownian motion on a tracial $W^{*}$-probability space $(\mathcal{A},\phi)$ with canonical filtration $(\mathcal{F}_{t})_{t\geq 0}$. For any $F\!\in\!L^{2}(\mathcal{A},\phi)$ adapted to $(\mathcal{F}_{t})$ whose free Malliavin derivative $D_{t}F$ belongs to $L^{2}(\mathcal{A}\,\bar{\otimes}\,\mathcal{A},\phi\!\otimes\!\phi)$, there exists a unique adapted process $u_{t}\in L^{2}_{\mathrm{ad}}([0,T];\mathcal{A})$ such that
$F=\phi[F]+\int_{0}^{T}\mathbb{E}_{\phi}[D_{t}F\mid\mathcal{F}_{t}]\,\#\,dS_{t},$ (9)
where #\# denotes the free stochastic integral and $\mathbb{E}_{\phi}[\cdot\mid\mathcal{F}_{t}]$ the conditional expectation in $(\mathcal{A},\phi)$.
The proof follows standard arguments in free Malliavin calculus and is given in Appendix C (see also [14, 35]).
@@ VI Reverse-Time Free SDE: Malliavin and PDE Derivations
extends Voiculescu’s classical free integration-by-parts identity from noncommutative polynomials to the full Sobolev space $W^{1,2}_{\mathrm{free}}(\mu_{t})$.
@@
Let $F\in\mathrm{Dom}(D)$ and approximate it by a sequence of self-adjoint noncommutative polynomials $\{P_{n}(X_{t})\}_{n\geq 1}$ such that $P_{n}(X_{t})\to F$ in $L^{2}(\mathcal{A},\tau)$ and $DP_{n}(X_{t})\to DF$ in $L^{2}(\mathcal{A}\otimes\mathcal{A})$. For each polynomial $P_{n}$, Voiculescu’s original integration-by-parts formula [12, Prop. 7.1] gives
@@
where $\mu_{X_{0}}$ is compactly supported with $\chi(\mu_{X_{0}})>$-$\infty$ and $\Phi^{\ast}(\mu_{X_{0}})<\infty$.
  1. (i)
    The forward equation (F) admits a unique strong solution $(X_{t})_{0\leq t\leq T}\subset L^{2}(\mathcal{A},\varphi)$ in the sense of free Itô calculus, and for each $t$ the spectral law $\mu_{X_{t}}$ possesses a conjugate variable $\Xi_{t}\in L^{2}(W^{\ast}(X_{t}),\varphi)$. Existence and uniqueness in law follow from Kargin [36] and Biane–Speicher [14].
  2. (ii)
    The time-reversed process $\widetilde{X}_{s}\mathrel{\mathop{\ordinarycolon}}=X_{T-s}$ satisfies the reverse-time free stochastic differential equation
@@
    where $(\widetilde{S}_{s})_{s\in[0,T]}$ is a backward free Brownian motion adapted to the reversed filtration $\widetilde{\mathcal{F}}_{s}=\sigma\{\widetilde{X}_{u}\mathrel{\mathop{\ordinarycolon}}0\leq u\leq s\}$. Equation (R) is well posed in $L^{2}(\mathcal{A},\varphi)$ and its solution is unique in law.
@@
Step 1: Existence and forward representation. Let $\Lambda(t)=\int_{0}^{t}\beta(r)\,\mathrm{d}r$. By the free Itô integration theory of Biane and Speicher [14] and the well-posedness results of Kargin [36], the SDE (F) admits a unique strong $L^{2}$ solution
$X_{t}=e^{-\frac{1}{2}\Lambda(t)}X_{0}+\!\int_{0}^{t}e^{-\frac{1}{2}(\Lambda(t)-\Lambda(u))}\sqrt{\beta(u)}\,\mathrm{d}S_{u},$
which satisfies $\|X_{t}\|_{2}<\infty$ for all $t\in[0,T]$ by the free Itô isometry.
Step 2: Clark–Ocone representation. For fixed $T$, the free Clark–Ocone formula (Lemma 1, proved in Appendix C) provides the decomposition
@@
Inserting the explicit forward solution and applying the free Itô isometry yields
@@
Since the increments $(\mathrm{d}S_{t})$ are freely independent and $u_{t}$ is $\mathcal{F}_{t}$-adapted, conditional expectation removes future stochastic terms, leaving the deterministic factor $e^{-\frac{1}{2}(\Lambda(T)-\Lambda(t))}$.
Step 3: Identification of the drift correction. By Lemma 3, for the log-density functional $G=\log\pi_{t}(X_{t})$ we have $\mathbb{E}_{\varphi}[D_{t}G\mid\mathcal{F}_{t}]=-\,\Xi_{t}$. Combining this with the forward drift term $\tfrac{1}{2}\beta(t)X_{t}$ in (F) produces the backward drift $\tfrac{1}{2}\beta(t)X_{t}-\beta(t)\Xi_{t}$.
Step 4: Time reversal and uniqueness in law. Define $\widetilde{X}_{s}=X_{T-s}$. By Biane’s time-reversal duality for free stochastic integrals [37, Prop. 3.1], the process $\widetilde{S}_{s}\mathrel{\mathop{\ordinarycolon}}=S_{T}-S_{T-s}$ is a backward free Brownian motion adapted to $\widetilde{\mathcal{F}}_{s}$. Applying free Itô calculus for reversed filtrations yields
$\mathrm{d}\widetilde{X}_{s}=\Bigl(-\tfrac{1}{2}\beta(T-s)\widetilde{X}_{s}-\beta(T-s)\,\Xi_{T-s}\Bigr)\mathrm{d}s+\sqrt{\beta(T-s)}\,\mathrm{d}\widetilde{S}_{s}.$
The square-integrability of $X_{t}$ and $\Xi_{t}$ implies integrability of all coefficients, ensuring existence and uniqueness in law of the solution to (R) in $L^{2}(\mathcal{A},\varphi)$ by Kargin’s general theory of free SDEs [36]. Hence $\widetilde{X}_{s}$ is the unique weak solution of (R), completing the proof. ∎
@@
Both derivations show that the term $\Xi_{s}$ plays the role of the noncommutative score function in the reverse-time dynamics, and that the reversed diffusion acts as a free deconvolution flow.